# Multiplayer Nash Preference Optimization

### v1

@@ Abstract
Reinforcement learning from human feedback (RLHF) has emerged as the standard paradigm for aligning large language models (LLMs) with human preferences. However, reward-based methods built on the Bradley–Terry assumption struggle to capture the non-transitive and heterogeneous nature of real-world preferences. To address this, recent studies have reframed alignment as a two-player Nash game, giving rise to Nash learning from human feedback (NLHF). While this perspective has inspired algorithms such as INPO, ONPO, and EGPO with strong theoretical and empirical guarantees, they remain fundamentally restricted to two-player interactions, creating a single-opponent bias that fails to capture the full complexity of realistic preference structures. In this work, we introduce Multiplayer Nash Preference Optimization (MNPO), a novel framework that generalizes NLHF to the multiplayer regime. It formulates alignment as an $n$-player game, where each policy competes against a population of opponents while being regularized toward a reference model. Our framework establishes well-defined Nash equilibria in multiplayer settings and extends the concept of duality gap to quantify approximation quality. We demonstrate that MNPO inherits the equilibrium guarantees of two-player methods while enabling richer competitive dynamics and improved coverage of diverse preference structures. Through comprehensive empirical evaluation, we show that MNPO consistently outperforms existing NLHF baselines on instruction-following benchmarks, achieving superior alignment quality under heterogeneous annotator conditions and mixed-policy evaluation scenarios. Together, these results establish MNPO as a principled and scalable framework for aligning LLMs with complex, non-transitive human preferences. Code is available at https://github.com/smiles724/MNPO.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress in instruction following and open-ended reasoning, primarily through reinforcement learning from human feedback (RLHF) (Christiano et al., 2017). In RLHF, annotators provide preference comparisons between model outputs, and these signals are used to align LLMs with human expectations. While reward-based RLHF pipelines built upon the Bradley-Terry (Bradley & Terry, 1952) model have enabled widely deployed systems (e.g., InstructGPT (Ouyang et al., 2022), Claude (Bai et al., 2022), and Gemini (Team et al., 2023)), they assume transitive preferences and scalar reward functions. Recent empirical studies in LLM alignment have revealed that human preferences often exhibit non-transitive patterns and heterogeneous structures that challenge these assumptions (Ethayarajh et al., 2024; Wu et al., 2024).
@@
This limitation has motivated recent attempts to develop game-theoretic formulations of alignment. Rather than relying on scalar rewards that enforce transitivity constraints, recent work treats preference optimization as finding Nash equilibria in games defined by general preference oracles (Munos et al., 2023). In this Nash learning from human feedback (NLHF) paradigm, a well-aligned policy represents an optimal strategy that cannot be exploited by competing policies. At equilibrium, policies achieve a balanced win rate that reflects strategic optimality rather than mediocrity. Subsequent studies have explored this paradigm through various algorithmic lenses. For instance, Zhang et al. (2025b) leverages no-regret learning to approximate the Nash equilibrium via self-play. Zhang et al. (2025a) accelerates convergence using optimistic mirror descent. Zhou et al. (2025) establishes last-iterate convergence with extragradient updates. These approaches advance both theoretical guarantees and empirical stability compared to traditional RLHF.
Despite advances, the existing NLHF remains constrained to two-player settings where a single policy competes against one opponent. However, real-world preference alignment often involves diverse annotators, heterogeneous evaluation criteria, or mixtures of historical model checkpoints - contexts that are better modeled as multiplayer games (Freund & Schapire, 1999). Extending NLHF to the multiplayer regime raises new challenges: (i) how to formulate objectives that balance fairness and competitiveness among multiple policies, (ii) how to design algorithms that converge to multiplayer Nash equilibria, and (iii) how to ensure tractability when scaling to larger LLMs (Sokota et al., 2022).
@@
We address these challenges by proposing Multiplayer Nash Preference Optimization (MNPO), a principled framework that generalizes two-player preference optimization to $n$-player games. In MNPO, each policy simultaneously competes against all other policies while being regularized toward a reference policy, creating a competitive equilibrium that balances performance against the population with adherence to a trusted baseline. Our contributions are threefold:
@@
  • Theoretical Framework: We establish that MNPO admits natural equilibrium characterizations, including well-defined Nash policies and duality gaps that measure alignment quality. We prove that MNPO inherits the desirable convergence properties of existing two-player formulations while enabling richer equilibrium dynamics.
  • Algorithmic Innovation: We introduce time-dependent MNPO (TD-MNPO), where opponent sets evolve adaptively using weighted combinations of historical policies. This approach enables models to dynamically incorporate past knowledge while maintaining training stability.
@@
Our analysis reveals that MNPO provides a unifying perspective on preference optimization, subsuming many existing methods as special cases while offering improved robustness in multi-agent alignment scenarios. By bridging recent advances in NLHF with the practical demands of aligning LLMs to diverse and potentially non-transitive human preferences, MNPO establishes a scalable foundation for next-generation alignment techniques.
@@ 2 RLHF Preliminaries
The prevalent RLHF framework (Christiano et al., 2017; Ouyang et al., 2022) assumes the Bradley-Terry model. It assumes that there exists a reward function $r^{*}$ such that for any $x\in\mathcal{X}$ and $y^{1},y^{2}\in\mathcal{Y}$, we have $\mathbb{P}\left(y^{1}\succ y^{2}\mid x\right)=\frac{\exp\left(r^{*}\left(x,y^{1}\right)\right)}{\exp\left(r^{*}\left(x,y^{1}\right)\right)+\exp\left(r^{*}\left(x,y^{2}\right)\right)}=\sigma\left(r^{*}\left(x,y^{1}\right)-r^{*}\left(x,y^{2}\right)\right).$
After learning a reward function $R(\cdot,\cdot)$, RLHF algorithms aim to maximize the following KL-regularized objective with preference optimization RL algorithms such as PPO (Schulman et al., 2017):
@@
The aforementioned algorithms rely on the Bradley-Terry model assumption, which may not always hold in practice. Recent studies (Munos et al., 2023; Calandriello et al., 2024; Ye et al., 2024; Zhang et al., 2025b; a) have instead directly considered the general preference distribution $\mathbb{P}$ without imposing additional assumptions, framing the preference optimization problem as a two-player game. These works assume the existence of a preference oracle $\mathbb{P}:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\rightarrow[0,1]$. It can be queried to obtain binary preference signals as $z\sim\operatorname{Ber}\left(\mathbb{P}\left(y^{1}\succ y^{2}\mid x\right)\right)$, where $z=1$ indicates that $y^{1}$ is preferred to $y^{2}$, and $z=0$ indicates the opposite. The preference distribution is introduced as $\lambda_{\mathbb{P}}\left(x,y,y^{\prime}\right)=\left(y,y^{\prime}\right)\cdot\mathbb{I}\left[U<\mathbb{P}\left(y\succ y^{\prime}\mid x\right)\right]+\left(y^{\prime},y\right)\cdot\mathbb{I}\left[U\geq\mathbb{P}\left(y\succ y^{\prime}\mid x\right)\right]$, where $U\sim\operatorname{Uniform}(0,1)$ is a random variable.
Given two policies $\pi_{1}$ and $\pi_{2}$, LLMs are aligned using this general preference oracle. Consequently, the game objective is written as:
@@ 3.1 RLHF as Multiplayer Games
To extend the two-player preference optimization objective to a multiplayer setting $\{\pi_{1},\pi_{2},...,\pi_{n}\}$, we consider a framework where each policy seeks to maximize its average preference probability against all other policies while regularizing toward a reference policy. We present the formulation of multiplayer games under two distinct assumptions: (i) the preference ranking assumption and (ii) the general preference oracle.
@@
The corresponding negative log-likelihood for a single comparison is:
$\log\mathbb{P}\left(y^{i}\succ\left\{y^{j}\right\}_{j\neq i}\Bigm|x\right)=\log\left(\exp\left(R\left(x,y^{i}\right)\right)+\sum_{j\neq i}\exp\left(R\left(x,y^{j}\right)\right)\right)-R\left(x,y^{i}\right).$ (6)
The generalized reward learning objective becomes:
[…]
@@
Multiplayer General Preference Optimization.
In addition to the Plackett-Luce assumption, we consider $n$-player preference optimization and leverage a universal preference oracle $\mathbb{P}:\mathcal{X}\times\mathcal{Y}\times\{\mathcal{Y}\}^{n-1}\rightarrow[0,1]$. It directly compares $y^{i}$ with a group of responses $\{y^{j}\}_{j\neq i}$, leading to the binary preference signals $z\sim\text{Ber}\left(\mathbb{P}\left(y^{i}\succ\left\{y^{j}\right\}_{j\neq i}\mid x\right)\right)$. Subsequently, each policy $\pi_{i}$ competes against other n1n-1 players, and the objective function becomes:
@@
Here, each policy $\pi_{i}$ seeks to maximize its expected preference over all other policies $\left\{\pi_{j}\right\}_{j\neq i}$. Meanwhile, the regularization term is kept to penalize $\pi_{i}$ with a KL divergence from the reference policy $\pi_{\text{ref}}$ weighted by $\tau$. This ensures that policies remain anchored to $\pi_{\text{ref}}$ to prevent overoptimization and maintain behavioral consistency. Regarding the dynamics of this multiplayer game, all policies $\left\{\pi_{i}\right\}_{i=1}^{n}$ are updated concurrently. Each policy $\pi_{i}$ optimizes its own $J$, leading to a competitive equilibrium where policies balance performance against the population and adherence to $\pi_{\mathrm{ref}}$.
Eq.8 has several key properties. (i) The symmetric treatment. All policies are treated equally, competing in a symmetric fashion. Therefore, $\pi_{1}^{*}=\pi_{2}^{*}=...=\pi_{n}^{*}$. (ii) Decentralized optimization. Each policy’s update depends only on its own actions and aggregate opponent behavior, avoiding complex interdependencies. (iii) Generalization of the two-player case. When $n=2$, each policy’s objective reduces to maximizing its pairwise preference probability minus its own KL penalty as in Eq. 2, aligning conceptually with competitive regularization.
@@
In this $n$-player game, the Nash equilibrium is a policy $\pi^{*}$ where no player can improve their objective by unilaterally deviating. Formally, for all $\pi_{i}\in\Pi$ :
@@
Due to symmetry, and assuming the KL terms vanish, the average win rate of the equilibrium policy $\pi^{*}$ against n1n-1 copies of itself is $\frac{1}{n}$. That is, $J\left(\pi^{*},\{\pi^{*}\}_{j\neq i}\right)=\frac{1}{n}$, analogous to the two-player case.
The duality gap quantifies how far a given policy $\pi$ is from the Nash policy $\pi^{*}$. Extending the two-player definition to an $n$-player setting, we define the set of opponent players as $O_{\pi}=\left\{\pi_{j}\right\}_{j=1}^{n-1}$. The duality gap in multiplayer games is then given by:
$\operatorname{DualGap}(\pi):=\mathbb{E}_{\pi_{j}\in O_{\pi}}\left[\max_{\pi_{j}}\min_{O_{\pi}\setminus\pi_{j}}J\left(\pi_{j},\pi\cup O_{\pi}\setminus\pi_{j}\right)\right]-\min_{O_{\pi}}J\left(\pi,O_{\pi}\right).$ (10)
This gap quantifies the maximum advantage a player could gain by unilaterally deviating from $\pi$ against the worst-case configuration of opponents, relative to the minimum payoff $\pi$ could incur if the opponents themselves deviate. Notably, the equilibrium condition is satisfied if and only if DualGap $\left(\pi^{*}\right)=0$, ensuring no player has an incentive to deviate unilaterally. Furthermore, if $\operatorname{DualGap}(\pi)\leq\epsilon$, then $\pi$ is defined as an $\epsilon$-approximate Nash policy in the multiplayer setting.
@@
There are well-known algorithms that approximately solve the Nash equilibrium in a constant-sum multiplayer game. In this work, we follow (Freund & Schapire, 1999) to establish an iterative framework that can asymptotically converge to the optimal policy on average. Given a learning rate $\eta$ of online mirror descent update, we start with a theoretical analysis that conceptually solves the multiplayer game as follows  [^1] For notational conciseness and avoiding confusion, we use a superscript to denote time here, so that $\pi^{(t)}$ is equivalent to $\pi_{t}$. :
$\pi_{i}^{(t+1)}(y\mid x)\propto\left(\prod_{j\neq i}\pi_{j}^{(t)}(y\mid x)\right)^{\frac{1}{n-1}}\exp\left(\frac{\eta}{n-1}\sum_{j\neq i}\mathbb{P}\left(y\succ\pi_{j}^{(t)}\mid x\right)\right)$ (11)
Eq. 11 is an iterative framework that relies on the multiplicative weight update and enjoys a clear structure. Initially, we have a base policy $\pi^{(0)}$, usually from some supervised fine-tuned model $\pi_{\text{ref}}$. In each iteration $t$, the updated policy $\pi^{(t+1)}$ is obtained from the reference policy $\pi^{(t)}$ following the multiplicative weight update. Particularly, a response $y$ should have a higher probability weight if it has a higher average advantage over the current policy $\pi^{(t)}$. Equivalently, Eq. 11 can be written as $\pi^{(t+1)}_{i}(y\mid x)={\Pi_{j\neq i}{\pi^{(t)}_{j}(y\mid x)}^{\frac{1}{n-1}}\exp\left(\frac{\eta}{n-1}\mathbb{P}\left(y\succ\pi^{(t)}_{j}\mid x\right)\right)}/{Z_{\pi^{(t)}}({x})}$, where $Z_{\pi^{(t)}}({x})$ is the normalization factor (a.k.a, the partition function). Then for any fixed $x$ and $y$, each ideal update policy $\pi^{(t+1)}_{i}$ should satisfy:
@@
The proof is deferred to Appendix E.1. Moreover, we replace the tricky term $\mathbb{P}\left(y\succ\pi^{(t)}_{j}\right)$ with a hyperparameter $\eta$ and propose the following loss to bypass it:
@@
See the proof in Appendix E.2. Here, the response pair $(y,y^{\prime})$ is directly sampled from the current policy $\pi^{(t)}$, which is crucial for the equivalence between $L_{t}^{\prime}(\pi)$ and $L_{t}(\pi)$.

### v2

@@ Abstract
Reinforcement learning from human feedback (RLHF) has emerged as the standard paradigm for aligning large language models with human preferences. However, reward-based methods built on the Bradley–Terry assumption struggle to capture the non-transitive and heterogeneous nature of real-world preferences. To address this, recent studies have reframed alignment as a two-player Nash game, giving rise to Nash learning from human feedback (NLHF). While this perspective has inspired algorithms such as INPO, ONPO, and EGPO with strong theoretical and empirical guarantees, they remain fundamentally restricted to two-player interactions, creating a single-opponent bias that fails to capture the full complexity of realistic preference structures. This work introduces Multiplayer Nash Preference Optimization (MNPO), a novel framework that generalizes NLHF to the multiplayer regime. It formulates alignment as an $n$-player game, where each policy competes against a population of opponents while being regularized toward a reference model. We demonstrate that MNPO inherits the equilibrium guarantees of two-player methods while enabling richer competitive dynamics and improved coverage of diverse preference structures. Comprehensive empirical evaluation shows that MNPO consistently outperforms existing NLHF baselines on instruction-following benchmarks, achieving superior alignment quality under heterogeneous annotator conditions and mixed-policy evaluation scenarios. Together, these results establish MNPO as a principled and scalable framework for aligning LLMs with complex, non-transitive human preferences. Code is available at https://github.com/smiles724/MNPO.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress in instruction following and open-ended reasoning through reinforcement learning from human feedback (RLHF) (Christiano et al., 2017). Traditional RLHF pipelines built upon the Bradley-Terry (Bradley and Terry, 1952) model have enabled widely deployed systems (e.g., InstructGPT (Ouyang et al., 2022), Claude (Bai et al., 2022), and Gemini (Team et al., 2023)), but assume transitive preferences and scalar reward functions. Recent empirical studies reveal that human preferences often exhibit non-transitive patterns and heterogeneous structures that violate these assumptions (Ethayarajh et al., 2024; Wu et al., 2024).
@@
This has motivated game-theoretic formulations of alignment that treat preference optimization as finding Nash equilibria in games defined by general preference oracles (Munos et al., 2023). In this Nash learning from human feedback (NLHF) paradigm, a well-aligned policy represents an optimal strategy that competing policies cannot exploit, achieving strategic optimality rather than mediocrity. Subsequent work has explored this paradigm through no-regret learning (Zhang et al., 2025b), optimistic mirror descent (Zhang et al., 2025a), and extragradient updates (Zhou et al., 2025), advancing both theoretical guarantees and empirical stability over traditional RLHF.
Despite these advances, existing NLHF formulations remain constrained to two-player settings, where a single policy competes against one opponent. However, real-world preference alignment rarely resembles a two-agent interaction. Instead, it often involves a mixture of annotators, heterogeneous evaluation criteria, multiple reward models, or even a sequence of historical model checkpoints—creating inherently multi-source, and sometimes conflicting, preference signals (Freund and Schapire, 1999). Collapsing this complex landscape into a single opponent introduces a bottleneck: the policy is optimized against only one distribution at a time, resulting in oscillatory behaviors, narrow exploration, and a brittle approximation of the broader preference population.
These limitations highlight the need for a framework that explicitly models alignment as competition against an entire population rather than a single synthetic opponent. In this work, we show that extending to the multiplayer setting (Freund and Schapire, 1999) provides a principled mean-field approximation that reduces gradient variance, stabilizes optimization, and more precisely captures diverse preference structures. Crucially, when all policies share the same preference oracle—as naturally occurs when competing against historical versions of a single policy trajectory—the resulting symmetric game admits strong theoretical guarantees via multiplicative weights update.
@@
We address these challenges by proposing Multiplayer Nash Preference Optimization (MNPO), a principled framework that generalizes two-player preference optimization to $n$-player games. In MNPO, each policy simultaneously competes against a population of other policies while being regularized toward a reference model, producing a competitive equilibrium that balances performance against the population with adherence to a trusted baseline. Our contributions are threefold:
@@
  • Theoretical Framework: We establish that MNPO with homogeneous preference oracles admits natural equilibrium characterizations, including well-defined Nash policies and duality gaps that measure alignment quality. We prove that MNPO inherits the desirable convergence properties of existing two-player formulations while enabling richer equilibrium dynamics.
  • Algorithmic Innovation: We introduce TD-MNPO, where opponent sets evolve adaptively using weighted combinations of historical policies, naturally yielding provable convergence guarantees. We further propose a heterogeneous extension (HT-MNPO) that, while lacking formal guarantees, demonstrates strong empirical performance across diverse preference sources.
@@
Our analysis reveals that MNPO provides a unifying perspective on RLHF, subsuming many existing methods as special cases while offering improved robustness in multi-agent alignment scenarios. By bridging recent advances in NLHF with the practical demands of aligning LLMs to diverse and potentially non-transitive human preferences, MNPO establishes a scalable foundation for next-generation alignment techniques.
@@ 2 RLHF Preliminaries
The prevalent RLHF framework (Christiano et al., 2017; Ouyang et al., 2022) assumes the Bradley-Terry model. It assumes that there exists a reward function $r^{*}$ such that for any $x\in\mathcal{X}$ and $y^{1},y^{2}\in\mathcal{Y}$, we have $\mathbb{P}\left(y^{1}\succ y^{2}\mid x\right)=\frac{\exp\left(r^{*}\left(x,y^{1}\right)\right)}{\exp\left(r^{*}\left(x,y^{1}\right)\right)+\exp\left(r^{*}\left(x,y^{2}\right)\right)}=\sigma\left(r^{*}\left(x,y^{1}\right)-r^{*}\left(x,y^{2}\right)\right).$ After learning a reward function $R(\cdot,\cdot)$, RLHF algorithms aim to maximize the following KL-regularized objective with preference optimization RL algorithms such as PPO (Schulman et al., 2017):
@@
The Bradley-Terry model assumption may not always hold in practice. Recent studies (Munos et al., 2023; Calandriello et al., 2024; Ye et al., 2024; Zhang et al., 2025b; a) have instead directly considered the general preference distribution $\mathbb{P}$ without imposing additional assumptions, framing the preference optimization problem as a two-player game. They assume the existence of a preference oracle $\mathbb{P}:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\rightarrow[0,1]$. It can be queried to obtain binary preference signals as $z\sim\operatorname{Ber}\left(\mathbb{P}\left(y^{1}\succ y^{2}\mid x\right)\right)$, where $z=1$ indicates that $y^{1}$ is preferred to $y^{2}$, and $z=0$ indicates the opposite. The preference distribution is introduced as $\lambda_{\mathbb{P}}\left(x,y,y^{\prime}\right)=\left(y,y^{\prime}\right)\cdot\mathbb{I}\left[U<\mathbb{P}\left(y\succ y^{\prime}\mid x\right)\right]+\left(y^{\prime},y\right)\cdot\mathbb{I}\left[U\geq\mathbb{P}\left(y\succ y^{\prime}\mid x\right)\right]$, where $U\sim\operatorname{Uniform}(0,1)$ is a random variable. Given two policies $\pi_{1}$ and $\pi_{2}$, LLMs are aligned using this general preference oracle, and the game objective is written as:
@@ 3.1 RLHF as Multiplayer Games
To extend the two-player preference optimization objective to a multiplayer setting $\{\pi_{i}\}_{i=1}^{n}$, we consider a framework where each policy seeks to maximize its average preference probability against all other policies while regularizing toward a reference policy. We first focus on the homogeneous multiplayer setting where all players share the same preference oracle. This symmetric structure provides the foundation for our theoretical guarantees.
@@
The corresponding negative log-likelihood for a single comparison is: $\log\mathbb{P}\left(y^{i}\succ\left\{y^{j}\right\}_{j\neq i}\Bigm|x\right)=\log\left(\exp\left(R\left(x,y^{i}\right)\right)+\sum_{j\neq i}\exp\left(R\left(x,y^{j}\right)\right)\right)-R\left(x,y^{i}\right)$. Consequently, the generalized reward learning objective for learning $\hat{R}$ becomes:
$\arg\max_{R\in\mathcal{R}}\mathbb{E}_{\left(x,\left\{y^{1:k}\right\}\right)\sim\mathcal{D}}\mathbb{E}_{y^{i}\in\left\{y^{1:k}\right\}}\underbrace{\left[R\left(x,y^{i}\right)-\log\left(\exp\left(R\left(x,y^{i}\right)\right)+\sum_{j\neq i}\exp\left(R\left(x,y^{j}\right)\right)\right)\right]}_{\text{Per-comparison log-likelihood }}.$ (6)
@@
Homogeneous Multiplayer Preference Oracle.
We consider the case where all $n$ players share the same universal preference oracle $\mathbb{P}:\mathcal{X}\times\mathcal{Y}\times\{\mathcal{Y}\}^{n-1}\rightarrow[0,1]$. It directly compares $y^{i}$ with a group of responses $\{y^{j}\}_{j\neq i}$ and output binary preference signals $z\sim\text{Ber}\left(\mathbb{P}\left(y^{i}\succ\left\{y^{j}\right\}_{j\neq i}\mid x\right)\right)$. Then, each policy $\pi_{i}$ competes against other n1n-1 players, and the objective function becomes:
@@
Here, each policy $\pi_{i}$ maximizes its expected preference against all other players $\left\{\pi_{j}\right\}_{j\neq i}$ while remaining close to the reference policy $\pi{\mathrm{ref}}$ via a KL penalty. The KL term, weighted by $\tau$, prevents over-optimization and ensures behavioral stability. All policies are updated concurrently: at each step, player $i$ improves its own objective $J$, leading the population toward a competitive equilibrium that balances performance against opponents and adherence to $\pi_{\mathrm{ref}}$.
Eq. 7 exhibits several important properties. (i) Symmetric treatment: All policies are treated equally and compete in a symmetric manner, ensuring that $\pi_{1}^{*}=\pi_{2}^{*}=\cdots=\pi_{n}^{*}$ at equilibrium. (ii) Decentralized optimization: Each policy’s update depends only on its own actions and the aggregate behavior of its opponents, thereby avoiding complex interdependencies. (iii) Generalization of the two-player case: When $n=2$, each policy’s objective reduces to maximizing its pairwise preference probability subject to its own KL penalty, as shown in Eq. 2.
@@
In this $n$-player game with homogeneous preference oracles, the Nash equilibrium is a policy $\pi^{*}$ where no player can improve their distinctive objectives by unilaterally deviating. Formally, for all $\pi_{i}\in\Pi$ :
@@
To quantify how far a given policy $\pi$ is from the Nash policy $\pi^{*}$, we define the unified duality gap in the multiplayer setting. For player $i$ with opponents fixed as $O_{\pi}=\{\pi_{j}\}_{j=1}^{n=1}$,
$\operatorname{DualGap}(\pi)=\max_{\pi^{\prime}\in\Pi}J(\pi^{\prime},\,O_{\pi})-J(\pi,\,O_{\pi}).$ (9)
This measures the maximum gain player $i$ could achieve by unilaterally switching to an alternative strategy while all opponents remain fixed. A Nash equilibrium is characterized by $\operatorname{DualGap}(\pi^{*})=0$, and $\operatorname{DualGap}(\pi)\leq\epsilon$ indicates that $\pi$ is an $\epsilon$-approximate Nash policy.
@@
There are well-known algorithms that approximately solve the Nash equilibrium in a constant-sum multiplayer game. In this work, we follow (Freund and Schapire, 1999) to establish an iterative framework that can asymptotically converge to the optimal policy on average. Given a learning rate $\eta$ of online mirror descent update, we start with a theoretical analysis that conceptually solves the homogeneous multiplayer game as follows  [^1] For notational conciseness and avoiding confusion, we use a superscript to denote time here, so that $\pi^{(t)}$ is equivalent to $\pi_{t}$. :
$\pi_{i}^{(t+1)}(y\mid x)\propto\left(\prod_{j\neq i}\pi_{j}^{(t)}(y\mid x)\right)^{\frac{1}{n-1}}\exp\left(\frac{\eta}{n-1}\sum_{j\neq i}\left(y\succ\pi_{j}^{(t)}\mid x\right)\right).$ (10)
This iterative framework starts from a base policy $\pi^{(0)}$ such as $\pi_{\text{ref}}$. In each iteration, the updated policy $\pi^{(t+1)}$ is obtained from the reference policy $\pi^{(t)}$ following the multiplicative weight update. Particularly, a response $y$ should have a higher probability weight if it has a higher average advantage over the current policy $\pi^{(t)}$. Notably, Eq. 10 provides convergence guarantees to a Nash equilibrium in this homogeneous circustance, ensuring that the average policy $\bar{\pi}^{(T)}=\frac{1}{T}\sum_{t=1}^{T}\pi^{(t)}$ converges to an $\epsilon$-approximate Nash equilibrium with $\epsilon=O(1/\sqrt{T})$ regret bound (Hart and Mas-Colell, 2000).
Eq. 10 is equivalent to $\pi^{(t+1)}_{i}(y\mid x)={\Pi_{j\neq i}{\pi^{(t)}_{j}(y\mid x)}^{\frac{1}{n-1}}\exp\left(\frac{\eta}{n-1}\left(y\succ\pi^{(t)}_{j}\mid x\right)\right)}/{Z_{\pi^{(t)}}({x})}$, where $Z_{\pi^{(t)}}({x})$ is the normalization factor (a.k.a, the partition function). Then, for any fixed $x$ and $y$, each ideal update policy $\pi^{(t+1)}_{i}$ should satisfy:
@@
The proof is deferred to Appendix F.1. Moreover, we replace the tricky term $\mathbb{P}\left(y\succ\pi^{(t)}_{j}\right)$ with a hyperparameter $\eta$ and propose the following loss to bypass it:
@@
See the proof in Appendix F.2. Here, the response pair $(y,y^{\prime})$ is directly sampled from the current policy $\pi^{(t)}$, which is crucial for the equivalence between $L_{t}^{\prime}(\pi)$ and $L_{t}(\pi)$.